OpenStudy (anonymous):

Assuming that none of the solutions are undefined, simplify each. 12a^5/6a plz help!!

OpenStudy (mertsj):

\[\frac{12a^5}{6a}=\frac{12}{6}\times\frac{a^5}{a}=2a^4\]

OpenStudy (anonymous):

Thank you

OpenStudy (mertsj):

yw